[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Tue Jul 19, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Wed Jul 20 02:57:59 UTC 2016


HSU Shift Log for Tue Jul 19 2016 22:57:54 GMT-0400 (EDT)

Shift: R

Forecaster: E Blake

Technical Issues Summary:

ATCF issues: Both satellite fixes imported without a problem tonight.

N-AWIPS/AWIPS2 issues: None

Other Techincal issues: None
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160720/50f350d1/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list